Aray of light passes from vacuum into a medium of refractive index m, the angle of incidence is found to be twice the angle of refraction. Then the angle of incidence is
step by sep solution please

Grade:12

1 Answers

Arun
5 years ago

Let angle of incidence equal to I.

So angle of refraction will be I/2.

sin I/(sin I/2)=m. (Snell's law)

sin I =nsin I/2

But sin I =2sin I/2cosi/2

After substituting and simplifying,

We get:-

cos I/2=m/2

I/2=cos^-1(m/2)

I=2cos^-1(m/2)